OpenStudy (anonymous):

PLEASE HELP I GIVE MEDALS! Simplify (2m^2 n^-3)^4

OpenStudy (solomonzelman):

\[(2m^2 n^{-3})^4\]this?

OpenStudy (solomonzelman):

\[(2m^{2 \times 4}n^{-3 \times 4})\]\[(2m^{8}n^{-12})\] being that \[a^{-b}=\frac{1}{a^b}~~~~~~~~~~~~~~~~there fore\]\[n^{-12}=\frac{1}{n^{12}}\]\[\frac{2m^8}{n^{12}}\]

OpenStudy (solomonzelman):

Makes sense?